# Rejoining fragmented ancient bamboo slips with physics-driven deep learning

### v1

@@ Discussion
The source code for WisePanda framework is publicly available on GitHub at https://github.com/zhujinchi/wisepanda. The repository contains the implementation of our physics-driven deep learning model for bamboo slip fragment matching, including the fracture modeling algorithms, corrosion simulation processes, neural network architecture and user interface. Users can install the required dependencies by following the installation instructions provided in the repository. The codebase is designed to be accessible to researchers, featuring an intuitive graphical user interface for fragment selection, comparison, and verification with integrated AI assistance. Additionally, we have established a project website at https://wisepanda.tech/ that provides demonstration videos, visual examples, and detailed documentation of the fragment rejoining and visualization tools discussed in this paper. This open-source release aims to facilitate further research in digital archaeology and enable the application of our methods to other types of archaeological fragment matching challenges.
@@
Figure E1: Typical fracture patterns of ancient bamboo slips. (a), The upper row displays bamboo samples with transverse fractures, the most common fracture mode (approximately 70%), characterized by irregular wavy breakage patterns resulting from stress propagating perpendicular to the bamboo fiber direction. These fractures often have limited text preservation at the fracture edges, with the fragmented characters providing insufficient information for reliable matching, thus making the morphological features of these edges the primary basis for rejoining. (b), The lower row exhibits longitudinal fractures, a less common but significant fracture mode, characterized by relatively smooth fracture edges that occur along the bamboo fiber direction. With these smoother fracture patterns, text matching becomes the primary criterion for correct rejoining, as the physical interlocking features are less pronounced. This diversity of fracture patterns and preservation states highlights the intrinsic complexity inherent in ancient bamboo slip rejoining.
@@
Figure E2: Physics-driven model of bamboo fracture propagation. a, Bamboo microstructure showing vertical fiber bundles (red blocks) and a typical fracture path (dotted black line) traversing across fibers. The black arrows indicate shear forces ($\sigma^{\infty}Z^{+}$) inducing Mode III fracture. b, Bamboo fracture modeling with stress field propagation between fibers. Points $P_{1}$, $P_{2}$, and $P_{3}$ represent positions on consecutive fibers $f_{i-1}$, $f_{i}$ and $f_{i+1}$, with angles $\theta_{i-1}$ and $\theta_{i}$ showing fracture directions of fibers $f_{i-1}$and $f_{i}$. Blue dotted circles illustrate the stress field emanating from point $P_{2}$, determining the probability distribution of potential fracture paths.
@@
Figure E3: Data distribution analysis and matching performance. a, t-SNE distributions of fragment features showing the relationship between synthetic data (blue) and real bamboo fragments (green) across four parameter configurations. Progressive convergence indicates effective parameter optimization in our model. b, Fracture matching quality matrices for 118 real fragment pairs. Deeper blue indicates higher similarity, with diagonal elements representing correct matches. The lower-right matrix shows optimal discriminative performance with clear contrast between prominent diagonal (blue) and off-diagonal elements (white), corresponding to the most effective parameter setting (lower-left t-SNE plot). This confirms that generating more realistic synthetic data directly improves fragment matching accuracy.
@@
Figure E4: Robustness analysis of fragment matching algorithms with increasing interference fragments. Performance comparison across six accuracy levels (Top-1 to Top-100) as candidate pool expands from 0 to 1114 fragments. WisePanda (orange bars) maintains relatively stable performance despite increasing interference, while Manual Random Search (dotted purple line) shows rapid accuracy decline. Other methods (DTW, SIS, FMM, and Cosine Similarity) also exhibit significant performance drops. At maximum interference, WisePanda maintains 50-65% accuracy at Top-50 and Top-100 levels, while competing methods fall below 35%, demonstrating physics-driven approach’s advantage in real archaeological scenarios with thousands of potential matches.
@@
Figure E5: Practical application of WisePanda. a, WisePanda software interface showing two successful matching cases: target bamboo slip fragments (left) with their corresponding matching results (right), demonstrating the system’s verification capabilities. b, Efficiency comparison between WisePanda-assisted and traditional manual matching across five test cases. The table records target and result fragments, system recommendation rank, and time measurements for both methods. WisePanda reduces average matching time from 266.54 to 11.52 seconds—approximately a 20-fold efficiency improvement. In three of five cases, the correct match received top ranking (rank 1), confirming the algorithm’s accuracy. These results validate the physics-driven approach’s practical value for large-scale bamboo slip fragment rejoining.

### v2

@@ Methods
The source code for WisePanda framework is publicly available on GitHub at https://github.com/zhujinchi/wisepanda under Apache License 2.0. The repository contains the implementation of our physics-driven deep learning model for bamboo slip fragment rejoining, including the fracture modeling algorithms, corrosion simulation processes, neural network architecture, and user interface. Users can install the required dependencies by following the installation instructions provided in the repository. The codebase is designed to be accessible to researchers, featuring an intuitive graphical user interface for fragment selection, comparison, and verification with integrated AI assistance. Additionally, we have established a project website at https://wisepanda.tech/ that provides demonstration videos, visual examples, and detailed documentation of the fragment rejoining and visualization tools discussed in this paper. This open-source release aims to facilitate further research in digital archaeology and enable the application of our methods to other types of archaeological fragment rejoining challenges.
[…]
[title=References - Methods]
@@
Figure E1: Typical fracture patterns of ancient bamboo slips. (a), The upper row shows bamboo samples with transverse fractures, the most common fracture mode (approximately 70%), characterized by irregular wavy breakage patterns. Most transverse fracture curves contain no text, with only occasional fragments showing partial characters, making morphological features the primary basis for rejoining. (b), The lower row exhibits longitudinal fractures, characterized by relatively smooth fracture curves along the fiber direction. These fragments typically preserve complete text, making textual continuity the primary criterion for matching. This diversity of fracture patterns and text preservation highlights the complexity of ancient bamboo slip rejoining.
@@
Figure E2: Physics-driven model of bamboo fracture propagation. a, Bamboo microstructure showing vertical fiber bundles (red blocks) and a possible fracture path (composed of dotted black segments) traversing across fibers. The black arrows indicate shear forces ($\sigma^{\infty}Z^{+}$) inducing Mode III fracture. b, Bamboo fracture modeling with stress field propagation between fibers. Points $P_{1}$, $P_{2}$, and $P_{3}$ represent left fracture height on consecutive fibers $f_{i-1}$, $f_{i}$ and $f_{i+1}$, with angles $\theta_{i-1}$ and $\theta_{i}$ showing fracture directions of fibers $f_{i-1}$and $f_{i}$. Blue dotted circles illustrate the stress field emanating from point $P_{2}$, determining the probability distribution of potential fracture paths.
@@
Figure E3: Data distribution analysis and matching performance. a, t-SNE distributions of fragment curve features showing the relationship between synthetic data (blue) and real bamboo fragments (green) across two parameter configurations. Convergence indicates effective parameter optimization in model. b, Fracture matching quality matrices for 118 real fragment pairs. Deeper blue indicates higher similarity, with diagonal elements representing correct matches. The lower-right matrix shows good discriminative performance with clear contrast between prominent diagonal (blue) and off-diagonal elements (white), corresponding to the most effective parameter setting (right of a).
@@
Figure E4: Robustness analysis of different fragment matching algorithms with increasing interference fragments. Performance comparison across six accuracy levels (Top-1 to Top-100) as candidate pool expands from 0 to 1000 fragments. WisePanda (orange bars) maintains relatively stable performance despite increasing interference, while Manual Random Search (dotted purple line) shows rapid accuracy decline. Other methods (DTW, SIS and FMM) show lower performance compared to WisePanda across different interference levels. At maximum interference, WisePanda maintains 50-65% accuracy at Top-50 and Top-100 levels, while competing methods achieve below 36% at Top-50, demonstrating the physics-driven approach’s advantage in real archaeological scenarios with thousands of potential matches.
@@
Figure E5: Practical application of WisePanda. a, WisePanda software interface showing two successful matching cases: target bamboo slip fragments with their corresponding matching results, demonstrating the system’s verification capabilities. b, Efficiency comparison between WisePanda-assisted and traditional manual matching across five test cases. The table records query (top fragment) and result fragments, system recommendation rank, and time measurements for both methods. WisePanda reduces average matching time from 326.49 to 15.03 seconds—approximately a 20-fold efficiency improvement. These results validate the physics-driven approach’s practical value for large-scale bamboo slip fragment rejoining.